Please show all work!2. You are given the following : Claim sizes follow a normal distribution with mean i and variance o2 = 47, 300. The hyp othe
Answer #1

Given

X N (10000, 47300)

then

X - 10000 V 47300 ~ N(0,1

9600 - 10000 P(X < 9600) = P(Z < V 47300

= P( z < -1.84 )

= 0.0329

Therefore , P value = 0.0329 is the probability of getting the observed value (9600) given that u= 10000

P value > 0.02

Fail to reject H0 at 0.02 significance level

P value < 0.05

Reject H0 at 0.05 significance level , but not at 0.02 level

Note : If P value < significance level , we reject H0 at that significance level.